# Leak Proof PDBBind: A Reorganized Dataset of Protein-Ligand Complexes for More Generalizable Binding Affinity Prediction

### v2

@@ Abstract
Many physics-based and machine-learned scoring functions (SFs) used to predict protein-ligand binding free energies have been trained on the PDBBind dataset. However, it is controversial as to whether new SFs are actually improving since the general, refined, and core datasets of PDBBind are cross-contaminated with proteins and ligands with high similarity, and hence they may not perform comparably well in binding prediction of new protein-ligand complexes. In this work we have carefully prepared a cleaned PDBBind data set of non-covalent binders that are split into training, validation, and test datasets to control for data leakage, defined as proteins and ligands with high sequence and structural similarity. The resulting leak-proof (LP)-PDBBind data is used to retrain four popular SFs: AutoDock Vina, Random Forest (RF)-Score, InteractionGraphNet (IGN), and DeepDTA, to better test their capabilities when applied to new protein-ligand complexes. In particular we have formulated a new independent data set, BDB2020+, by matching high quality binding free energies from BindingDB with co-crystalized ligand-protein complexes from the PDB that have been deposited since 2020. Based on all the benchmark results, the retrained models using LP-PDBBind consistently perform better, with IGN especially being recommended for scoring and ranking applications for new protein-ligand systems.
@@ 1 Introduction
Scoring functions (SFs) are crucial in computer aided drug discovery, utilized for selecting the most probable binding pose geometry and/or free energy of binding between a ligand and a protein.1, 2 There are a plethora of SFs being developed and widely used by computational chemists, but they can be broadly categorized into either classical scoring functions (CSFs)3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, or machine learning scoring functions (MLSFs).14, 15, 16, 17, 18, 19, 20, 21 CSFs such as the commonly used AutoDock Vina3 are designed to model intermolecular interactions or missing free energy components, and can benefit both from better design of the functional form, many-body physics, as well as the availability of training data for parameterization of their semi-empirical or empirical functions.4, 3, 5, 6, 9, 7, 8. Data-driven MLSFs by contrast are less reliant on physical interaction modeling and are far more dependent on experimental information22, 23, 11, culminating in the current development and use of sophisticated machine learning (ML) models14, 15, 16, 17, 18, 19, 20, 21 whose much larger parameter spaces are optimized on large, high quality datasets. Some representative MLSF designs include Random Forest Score (RF-Score)14 that is based on atom-pair distance counts, InteractionGraphNet (IGN)15 that uses graph neural network (GNN) to represent raw 3-dimensional protein and ligand structures, and DeepDTA18 that predicts binding affinities using amino acid sequences of the proteins and SMILES strings of the ligands. More recent models have been proposed that achieve better accuracy at binding affinity prediction, including RFScore-V3 24 that incorporates terms from AutoDock Vina, and also PIGNet16, RTMScore17 and GIGN21 which are more advanced GNN-based models.
@@
The majority of protein-ligand interaction SF predictors, whether physical or machine-learned, have been trained on the PDBBind dataset.25 The Comparative Assessment of Scoring Functions (CASF) benchmark, which assesses the scoring power, ranking power, docking power and screening power of various SFs, was also conducted on the PDBBind core set.26 More specifically PDBBind is a curated set of $\sim$20K protein-ligand complex structures and their experimentally measured binding affinities, in which the ”general” and ”refined” data subsets are used for training, and a separate ”core” set is used for testing. However, not all protein-ligand complex structures in the general and refined data sets are as high-quality as those in the core set which contains protein-ligand complexes with the best structural resolution and most reliable binding affinity data.25 Additionally, the average size of the ligands in the core set is also smaller than that in the rest of the PDBBind dataset, which may make the core set easier prediction targets. Furthermore, as we will show later, the majority of the core data records in the PDBBind dataset have identical proteins and/or ligands with that found in the general and/or refined sets. As such, most empirical SFs and MLSF models have been trained with significant data leakage, and thus their reported performance on the core set is only a true measure of new protein-ligand complexes with high similarity, but will inevitably have limited transferability to low-similarity ligand-proteins scenarios.
@@
The definition of data leakage in PDBBind is not as literal a scenario in which the exact occurrence of input representation occurs in both training and test sets. Instead data leakage here refers to highly similar protein or ligands measured by sequence and chemical similarity and similar structure occurring in both the training and testing datasets. This will lead to artificially high evaluation results that do not fully reflect the performance of the models in real world applications.27 Therefore, the most standard data splitting protocol in the machine learning community, which is to randomly split the dataset into train, validation and test subsets will inevitably introduce significant data leakage issue. Given that many MLSFs training on PDBBind do not carefully investigate the data similarity issue, whether these new MLSFs have truly surpassed traditional CSFs in actual predictive performance is still an open question, or how well any given CSF or MLSF performs on new protein-ligand binding applications.28, 29, 30, 31 It is already known that any SF model can be overtrained so that it has exceptional performance on the training dataset, a problem which can be mitigated by several known regularization strategies to provide better generalization to an independent test set.32, 33 Equally important, however, is to control for data leakage into the test set itself, without which can lead to false confidence in predictive capacity when the training dataset has high similarity to the test set, but manifests as poor generalization when the sequence or structure similarity is low.29
@@
Two typical solutions have been taken to mitigate this data leakage issue: one is to perform data splits based on scaffolds, protein families or protein clusters, and the other is to perform time-based cutoff which was proposed on earlier versions of PDBBind24, and was suggested by the authors of the EquiBind model on the PDBBind 2020 release34. The scaffold or target based splitting can make sure the proteins in the complex are not similar between training and testing datasets, but they usually do not take the ligand similarities into account. The idea of a time-based splitting mimics a ”blind test” setting, by which the model can only be trained with data released in year 2019 or earlier, and predictions are made with data after year 2019, so that the test data will never be covered by training data. However, since new drugs are being developed that can interact with popular protein targets that have been established for years,35, 36 and existing drug molecules may also be tested on new proteins37, there are still frequent encounters with almost identical proteins or ligands in the latest experiments with earlier assays. As such, a time based splitting of the dataset is still not an ideal solution.
@@
For the purpose of retraining SFs that can either truly generalize, or just as importantly to anticipate when it will not work well when similarity of protein-ligand complexes is low, we build a new data split to mitigate data leakage of highly similar sequence and structures of both proteins and ligands in the training and test set. In this work we reorganize the PDBBind data into new train, validation and test datasets which we call Leak Proof PDBBind (LP-PDBBind), by minimizing sequence and chemical similarity of the proteins and ligands between the datasets as well as making sure the protein-ligand structural interaction patterns are different among the data splits as well. Our new data splitting can be regarded as an improvement over just a protein scaffold split by providing similarity control on the ligands as well, which we hypothesize will decrease the chance of a model making predictions by memorizing ligand structures.38, 39 Furthermore we have cleaned the PDBBind data to eliminate covalent bound ligand-protein complexes (thus focusing on non-covalent binding), certain ligands with very low frequency occurrences of certain atomic elements, to remove obvious steric clashes, as well as maintaining consistency in reported binding free energies and their units.
@@
We then use the cleaned LP-PDBBind data for the development of new versions of several foundational SFs, including AutoDock Vina3, RF-Score14, IGN15, and DeepDTA18. This provides a standardized way to train and benchmark different SFs with the exact same train, validation and test data. Furthermore, in order to provide a true independent benchmark for the new SFs that result from retraining using the cleaned LP-PDBBind, we created a new evaluation dataset, BDB2020+. BDB2020+ is compiled based on entries into the BindingDB dataset40 that were deposited after 2020, and further filtered according to the same similarity control criteria used for the development of the new LP-PDBBind. As a further test of ranking power, we additionally prepared two sets of experimental binding affinity data for different ligand complexes of the SARS-CoV-2 main protease (Mpro)41 and epidermal grow factor receptor (EGFR)42, neither of which was included in the LP-PDBBind training dataset (although Mpro has similar SARS-CoV-1 protease proteins in PDBBind). Finally, we perform two control evaluated the performance of retrained SFs on docked structures using the old Vina SF to mimic a more typical drug discovery scenario.
@@
The results show that all SFs trained on PDBBind show better performance on the LP-PDBBind test set due to data leakage, while the retrained models using the LP-PDBBind split are found to give a more true measure of performance. Furthermore, the newly trained SFs in general now perform much better on the completely independent BDB2020+ data set, showing that the new split can help with better generalization. Furthermore, even though the models were trained using cocrystal structures, improvements are also seen on redocked structures, thus demonstrating the applicability for the retrained SFs in actual practice. Hence, we believe LP-PDBBind provides a better way of utilizing the existing PDBBind dataset, and provides a more realistic and meaningful benchmark to help develop higher quality and more generalizable SFs in the future.
@@ 2 Methods
2.1 Cleaning the PDBBind dataset
The majority of protein-ligand affinity prediction models are designed for non-peptide drug-like small molecules. However, a careful investigation of the protein-ligand complexes in PDBBind dataset elucidated that all the ligands are not necessarily drug like. Figure 1(a-c) shows the distributions of the number of heavy atoms, molecular weights and quantitative estimate of drug-likeness (QED)43 for ligands in the general, refined and core set in the PDBBind dataset. Figure 1(d) shows that some ligands contain unusually large macrocycles, are peptide like, or contain long aliphatic chains. Including these data in the training of a ML model might make it harder for the model to learn key features related to the binding of drug-like small molecules to protein targets.
Some proteins and ligands in the PDBBind dataset contain uncommon elements, such as Hg, Cu and Sr in proteins or Co, Se or V in ligands; the proportion of occurrence for each element is summarized in Supplementary Table 3 for proteins and Supplementary Table 4 for ligands. A small portion of the structures in PDBBind dataset also contained steric clashes, and there was one structure for which the ligand was far away from any peptide chain (PDB ID: 2R1W). These should be excluded from the dataset due to the low quality of the complex structures, and the PDB IDs containing steric clashes (minimum heavy atom distance $<$1.75 Å) are also provided in the Supplementary Information. It is difficult for ML models to learn protein-ligand interactions when poor data and uncommon elements are present in the PDBBind dataset. Therefore, we have defined a ”cleaner” version of PDBBind which only contains all data with ligand QED values larger than 0.2, protein and ligand elements that have at least 1‰ (more than 19 occurrences in the dataset), and minimum heavy-atom distance between protein and ligand fall within 1.75 and 4 Å. We call this subset of the data Clean Level 1 (CL1).
Figure 1: Property distributions and example ligands in the PDBBind dataset. (a) Distribution of number of heavy atoms, (b) molecular weights and (c) QED values for ligands in general, refined and core set of PDBBind dataset. (d) Example of ligand structures in the PDBBind dataset, with the corresponding PDB IDs
The binding affinity in terms of $\Delta G$ is directly related to the dissociation coefficient $K_{d}$ or $K_{i}$ through the formula $\Delta G=-RT\ln(K)$.44 However, a large portion of the data in PDBBind is reported in terms of $\mathrm{IC}_{50}$, which cannot be easily translated to $\Delta G$s due to its dependence on other experimental conditions and inhibition mechanisms.45 The $\mathrm{IC}_{50}$ values for the same protein-ligand complex can vary up to one order of magnitude in different assays. In addition, some data in PDBBind were not reported as exact values. Therefore, a second clean level (CL2) was defined on top of CL1, that additionally requires the target values are converted from records with “$K_{d}=\mathrm{xxx}$” or “$K_{i}=\mathrm{xxx}$”, to ensure the reliability of experimental binding free energy data. Finally, considering the original splitting of PDBBind into general/refined/core was based on structure quality25, we also defined a third and highest quality level of data (CL3) that only retained data from refined and core set of PDBBind while still complying with all other quality-control metrics that we have defined. The total amount of data in PDBBind v2020 filtered with CL1, CL2 and CL3 are 14324, 7985 and 4404, respectively. The CL1 level dataset was used for retraining models described later because it contains the most amount of data. The CL2 level dataset was used for validation and test because the binding affinity data are more reliable at this level. We did not use CL3 in this study because it had the smallest amount of data, but as it may be useful for others we have provided identifiers of these protein-ligand complexes which are available at https://github.com/THGLab/LP-PDBBind/ along with the CL1, CL2 data sets.
@@ 2.2 New splitting of PDBBind dataset
We have formulated a new splitting of the PDBBind dataset to minimize the similarity between training, validation and test data as much as possible in order to eliminate the risk of data leakage. Ligand similarities were calculated as the Dice similarity46 between a pair of ligands based on Morgan fingerprints of the ligands using 1024 bits47. Similarities for proteins were calculated as the percentage of matched number of residues over the length of aligned sequences after sequences were aligned using the Needleman-Wunsch alignment48. More details on the calculation of similarity for proteins and ligands and data splitting procedure are given in the Supplementary Information. The total amount of data in the training, validation and test data after the new splitting are 11513, 2422 and 4860, respectively, and the data for each category is reported in Supplementary Table 5. The new splitting procedure ensures the data in the training set has a maximum protein sequence similarity of 0.5 and maximum ligand similarity of 0.99 to any data in the validation or test datasets, and the maximum similarity between any validation and test data is 0.9 for protein sequence similarity and 0.99 for ligand similarity. The similarity cutoffs are carefully chosen to reach a balance between the dis-similarity of train and test datasets and the amount of data in each subset. A higher similarity cutoff for the ligands was used compared to the protein similarity cutoff because similar ligands might interact differently with different types of proteins, which is useful for a ML model to learn.
Additionally, we explored the protein-ligand interaction similarities using proteo-chemometrics interaction fingerprints49 and validated the exclusion of highly similar interaction patterns in the dataset using our splitting procedure. The proteo-chemometric interaction fingerprints extend the connectivity fingerprints (ECFPs) by the interactions between ligand atoms and nearby residues from the protein in 3D space, and maps the interaction patterns into a fixed size integer vector with a length of 256. The pairwise interaction fingerprint similarities were then calculated using a weighted Jaccard similarity score
@@ 2.3 Compilation of the BDB2020+ dataset
Because many of the recent SFs have been trained on PDBBind, which means utilizing a subset of PDBBind as the test dataset risks data leakage, we have created a new benchmark dataset that is independent of PDBBind. To fulfill the need for a fair benchmark, we looked for the records deposited in BindingDB40, 50 after year 2020. Given that many data in BindingDB overlap with the PDBBind dataset which contains data up to year 2020, we have only looked at data in BindingDB deposited after year 2020 to minimize the chance of data leakage. BindingDB is one of the largest public binding affinity repositories, which also provides additional experimental conditions including assay information, pH and temperature. However, it does not guarantee each record has an associated 3D complex structure.
We have developed a workflow to match complex structures in RCSB PDB with records in BindingDB. Since a complex may contain multiple ligands, but only one matches the record in BindingDB, we selected the ligand in the PDB that has the best structural match with the SMILEs provided in BindingDB, and ensuring that the number of heavy atoms is exactly the same. We then used rdkit51 to reassign bond orders to the extracted ligands using the BindingDB SMILES as reference. This step was necessary because bond orders are usually not present in a PDB file and are typically inferred from local atomic geometries, which sometimes result in unreasonable bonding structures; thus the bond order reassignment step ensures the rationality of the processed structures.
@@
We have further prepared the redocked structures using the original AutoDock Vina for each protein-ligand complex in the BDB2020+ dataset to evaluate the performance of the retrained SFs on docked poses instead of co-crystal structures obtained from experiments, because the co-crystal structures are usually unavailable in real-world applications. Specifically, for each data in BDB2020+ dataset, we run AutoDock Vina docking to generate 30 poses with exhausiveness=64. The pose with the lowest Vina score and the pose that is most similar to the cocrystal structure in terms of symmetry-corrected RMSD 52 were both retained. However, any data with the lowest RMSD to the co-crystal structure larger than 2 Å were excluded to prevent the influence of bad docking, resulting a total of 104 data remained for the redocked set of BDB2020+ that is closest to the co-crystal structure.

### v3

@@ Abstract
The majority of machine learning scoring functions used in drug discovery for predicting protein-ligand binding poses and affinities have been trained on the PDBBind dataset. However, it is unclear whether these new scoring functions are actually an improvement over traditional models since often the training and test sets are cross-contaminated with proteins and ligands with high similarity, and hence they may not perform comparably well in binding prediction of unrelated protein-ligand complexes. In this work we have carefully prepared a new split of the PDBBind data set to control for data leakage, defined as proteins and ligands with high sequence and structural similarity. The resulting leak-proof (LP)-PDBBind data is used to retrain four popular SFs: AutoDock Vina, Random Forest (RF)-Score, InteractionGraphNet (IGN), and DeepDTA, to better test their capabilities when applied to new protein-ligand complexes. In particular we have formulated a new independent data set, BDB2020+, by matching high quality binding free energies from BindingDB with co-crystalized ligand-protein complexes from the PDB that have been deposited since 2020. Based on all the benchmark results, the retrained models using LP-PDBBind consistently perform better, with IGN especially being recommended for scoring and ranking applications for new protein-ligand systems.
@@ 1 Introduction
Many physics-based and machine-learned scoring functions (SFs) used to predict protein-ligand binding energies have been trained on the PDBBind dataset.40 The PDBBind dataset is a curated set of $\sim$20K protein-ligand complex structures and their experimentally measured binding affinities, in which the ”general” and ”refined” data subsets are typically used for training and/or validation, and a separate ”core” set is used for testing, in the development of new SFs. Significantly, however, the majority of the core data records in the PDBBind dataset have identical proteins and/or ligands with that found in the general and/or refined sets. As such, many empirical SF models have been trained with significant data leakage, and thus their reported performance on the core set is only a true measure of new protein-ligand complexes with high similarity, but will inevitably have limited transferability to low-similarity ligand-proteins scenarios.
@@
The definition of data leakage is not as literal a scenario in which the exact occurrence of input representation occurs in both training and test sets. Instead data leakage here refers to highly similar protein or ligands measured by sequence and chemical similarity and similar structure occurring in both the training and testing datasets. This will lead to artificially high evaluation results that do not fully reflect the performance of the models in real world applications for more diverse protein-ligand complexes.84 Therefore, the most standard data splitting protocol in the machine learning community, which is to randomly split the dataset into train, validation and test subsets will inevitably introduce significant data leakage issue. Given that many SFs training on PDBBind do not carefully investigate the data similarity issue, whether these new SFs have truly surpassed traditional SFs in actual predictive performance is still an open question, or how well any given SF performs on new protein-ligand binding applications.60, 38, 35, 79 It is already known that any SF model can be overtrained so that it has exceptional performance on the training dataset, a problem which can be mitigated by several known regularization strategies to provide better generalization to an independent test set.22, 71 Equally important, however, is to control for data leakage into the test set itself, without which can lead to false confidence in predictive capacity when the training dataset has high similarity to the test set, but manifests as poor generalization when the sequence or structure similarity is low.38
@@
Two typical solutions have been taken to mitigate this data leakage issue: one is to perform data splits based on scaffolds, protein families or protein clusters, and the other is to perform time-based cutoff which was proposed on earlier versions of PDBBind34, and was suggested by the authors of the EquiBind model on the PDBBind 2020 release65. The scaffold or target based splitting can make sure the proteins in the complex are not similar between training and testing datasets, but they usually do not take the ligand similarities into account. The idea of a time-based splitting mimics a ”blind test” setting, by which the model can only be trained with data released in year 2019 or earlier, and predictions are made with data after year 2019, so that the test data will never be covered by training data. However, since new drugs are being developed that can interact with popular protein targets that have been established for years,82, 49 and existing drug molecules may also be tested on new proteins57, there are still frequent encounters with almost identical proteins or ligands in the latest experiments with earlier assays. As such, a time based splitting of the dataset is still not an ideal solution. We note that other data-splitting methods29, 19 have been reported in 2025 without attribution to our work on the Arxiv posted in 2023.37
@@
For the purpose of retraining SFs that can either truly generalize, or just as importantly to anticipate when it will not work well when similarity of protein-ligand complexes is low, we build a new data split to mitigate data leakage of highly similar sequence and structures of both proteins and ligands in the training and test set. In this work we reorganize the PDBBind data into new train, validation and test datasets which we call Leak Proof PDBBind (LP-PDBBind), by minimizing sequence and chemical similarity of the proteins and ligands between the datasets as well as making sure the protein-ligand structural interaction patterns are different among the data splits as well. Our new data splitting can be regarded as an improvement over just a protein scaffold split by providing similarity control on the ligands as well, which we hypothesize will decrease the chance of a model making predictions by memorizing ligand structures.18, 69
@@
We use the new split for the development of new versions of several SFs, including AutoDock Vina67, RF-Score2, IGN28, and DeepDTA50. We note that DataSail29 and CleanSplit19, both reported in 2025, also retrained ML models but without attribution to our 2023 Arxiv paper.37 As we showed before, this provides a standardized way to train and benchmark different SFs with the exact same train, validation and test data. Furthermore, in order to provide a true independent benchmark for the new SFs that result from retraining using the new LP-PDBBind split, we created a new evaluation dataset, BDB2020+. BDB2020+ is compiled based on entries into the BindingDB dataset39 that were deposited after 2020, and further filtered according to the same similarity control criteria used for the development of the new LP-PDBBind. We note that DataSail29 also used our BDB2020+ data, which we generously provided to the community in the THG Lab github, but they used it without attribution.37 As a further test of ranking power, we additionally prepared two sets of experimental binding affinity data for different ligand complexes of the SARS-CoV-2 main protease (Mpro)13 and epidermal grow factor receptor (EGFR)25, neither of which was included in the LP-PDBBind training dataset (although Mpro has similar SARS-CoV-1 protease proteins in PDBBind). Finally, we perform two control evaluated the performance of retrained SFs on docked structures using the old Vina SF to mimic a more typical drug discovery scenario.
@@
The results show that all SFs trained on PDBBind show better performance on the test set due to data leakage, while the retrained models using the LP-PDBBind split are found to give a more true measure of performance. Furthermore, the newly trained SFs in general now perform much better on the completely independent BDB2020+ data set, showing that the new split can help with better generalization. Furthermore, even though the models were trained using co-crystal structures, improvements are also seen on redocked structures, thus demonstrating the applicability for the retrained SFs in actual practice. Hence, we believe the LP-PDBBind split provides a better way of utilizing the existing PDBBind dataset, and provides a more realistic and meaningful benchmark to help develop higher quality and more generalizable SFs in the future.
@@ 2 Methods
2.1 Data cleaning of PDBBind v2020
We have cleaned the PDBBind v2020 data to remove covalent ligand-protein complexes, the low populations of drug molecules with underrepresented chemical elements, and complexes with steric clashes; this is defined as clean level 1 (CL1). In addition to CL1, the CL2 level of cleaning aimed for consistent measures of binding free energies by either converting $K_{d}$ or eliminating data reported as $\mathrm{IC_{50}}$. All of the results reported here were based on training on CL1 data and testing on CL2. A more recent and thorough cleaning procedure, the HiQ-Bind workflow, has also been applied to PDBBind v2020 in that study.73
It is important to treat covalent and non-covalent binders separately, because most existing algorithms that predict protein-ligand binding primarily focus on non-covalent interactions. As far as we are aware, there has been no systematic study of whether the binders in PDBBind are covalent or non-covalent. Relying on the CovBinderInPDB 21 repository, we have identified covalent binders in the PDBBind dataset. The ligand names were extracted from the PDB files by comparing the minimum distance between any atom from the ligands in PDBBind database and residues in the PDB files downloaded from the RCSB database. If the minimum distance is less than 1Å, the matched residue name was compared with the record in CovBinderInPDB to identify whether the ligand is a covalent binder. If the minimum distance is more than 1Å or the residue name did not match the record in CovBinderInPDB, the structures were manually checked to identify whether the ligands are covalent binders or not. Ultimately 893 covalent binders and 18550 non-covalent binders were identified in the PDBBind dataset (see Supplementary Information). We also eliminated protein-ligand complexes with steric clashes, defined as a heavy-atom distance between protein and ligand that is below 1.75 Å (see Supplementary Information). In total there are 11,513 protein-ligand complexes in the training set, 2,422 in the validation set, and 4,860 in the test set before merging to create the new split (see Supplementary Table S1).
@@ 2.2 New splitting of PDBBind dataset
We have formulated a new splitting of the PDBBind dataset to minimize the similarity between training, validation and test data as much as possible in order to eliminate the risk of data leakage. Ligand similarities were calculated as the Dice similarity9 between a pair of ligands based on Morgan fingerprints of the ligands using 1024 bits59. Similarities for proteins were calculated as the percentage of matched number of residues over the length of aligned sequences after sequences were aligned using the Needleman-Wunsch alignment47. More details on the calculation of similarity for proteins and ligands and data splitting procedure are given in the Supplementary Information. The total amount of protein-ligand data in the training, validation and test data after the new splitting are 11513, 2422 and 4860, respectively, and the data for each category is reported in Supplementary Table S1. The new splitting procedure ensures the data in the training set has a maximum protein sequence similarity of 0.5 and maximum ligand similarity of 0.99 to any data in the validation or test datasets, and the maximum similarity between any validation and test data is 0.9 for protein sequence similarity and 0.99 for ligand similarity. The similarity cutoffs are carefully chosen to reach a balance between the dis-similarity of train and test datasets and the amount of data in each subset. A higher similarity cutoff for the ligands was used compared to the protein similarity cutoff because similar ligands might interact differently with different types of proteins, which is useful for a ML model to learn.
Additionally, we explored the protein-ligand interaction similarities using proteo-chemometrics interaction fingerprints70 and validated the exclusion of highly similar interaction patterns in the dataset using our splitting procedure. The proteo-chemometric interaction fingerprints extend the connectivity fingerprints (ECFPs) by the interactions between ligand atoms and nearby residues from the protein in 3D space, and maps the interaction patterns into a fixed size integer vector with a length of 256. The pairwise interaction fingerprint similarities were then calculated using a weighted Jaccard similarity score
@@
2.3 Retrained Models using the new split
The primary deliverable from this study are new SFs for protein-ligand binding prediction. We have considered both classical SFs as well as machine learned SFs. The commonly used AutoDock Vina67 is designed to model intermolecular interactions or missing free energy components, but could benefit from better training data for parameterization of their semi-empirical or empirical functions.45, 67, 12, 11, 72, 58, 4. Data-driven MLSFs by contrast are less reliant on physical interaction modeling and are far more dependent on experimental information17, 43, 46, culminating in the current development and use of sophisticated machine learning (ML) models2, 28, 44, 61, 50, 63, 42, 77 whose much larger parameter spaces are optimized on large, high quality datasets. Some representative MLSF models we consider include Random Forest Score (RF-Score)2 that is based on atom-pair distance counts, InteractionGraphNet (IGN)28 that uses graph neural network (GNN) to represent raw 3-dimensional protein and ligand structures, and DeepDTA50 that predicts binding affinities using amino acid sequences of the proteins and SMILES strings of the ligands. More recent models have been proposed that achieve better accuracy at binding affinity prediction, including RFScore-V3 34 that incorporates terms from AutoDock Vina, and also PIGNet44, RTMScore61 and GIGN77 which are more advanced GNN-based models.
@@ 2.4 Compilation of the BDB2020+ dataset
Because many of the recent SFs have been trained on PDBBind, which means utilizing a subset of PDBBind as the test dataset risks data leakage, we created a new benchmark dataset that is independent of PDBBind. To fulfill the need for a fair benchmark, we looked for the records deposited in BindingDB39, 15 after year 2020 until 2023. Given that many data in BindingDB overlap with the PDBBind dataset which contains data up to year 2020, we have only looked at data in BindingDB deposited after year 2020 to minimize the chance of data leakage. BindingDB is one of the largest public binding affinity repositories, which also provides additional experimental conditions including assay information, pH and temperature. However, it does not guarantee each record has an associated 3D complex structure.
We have developed a workflow to match complex structures in RCSB PDB with records in BindingDB. Since a complex may contain multiple ligands, but only one matches the record in BindingDB, we selected the ligand in the PDB that has the best structural match with the SMILEs provided in BindingDB, and ensuring that the number of heavy atoms is exactly the same. We then used rdkit33 to reassign bond orders to the extracted ligands using the BindingDB SMILES as reference. This step was necessary because bond orders are usually not present in a PDB file and are typically inferred from local atomic geometries, which sometimes result in unreasonable bonding structures; thus the bond order reassignment step ensures the rationality of the processed structures.
@@
We have further prepared the redocked structures using the original AutoDock Vina for each protein-ligand complex in the BDB2020+ dataset to evaluate the performance of the retrained SFs on docked poses instead of co-crystal structures obtained from experiments, because the co-crystal structures are usually unavailable in real-world applications. Specifically, for each data in BDB2020+ dataset, we run AutoDock Vina docking to generate 30 poses with exhausiveness=64. The pose with the lowest Vina score and the pose that is most similar to the cocrystal structure in terms of symmetry-corrected RMSD 36 were both retained. However, any data with the lowest RMSD to the co-crystal structure larger than 2 Å were excluded to prevent the influence of bad docking, resulting a total of 104 data remained for the redocked set of BDB2020+ that is closest to the co-crystal structure.